OpenStudy (anonymous):

Find the value of y so that the line passing through (6, y) and (7, 5) has a slope of -3.

OpenStudy (anonymous):

y = -3 x + b 5 = -21 + b b = 26 y = -3 x + 26 y = -18 + 26 y = 8

OpenStudy (karatechopper):

y=8

OpenStudy (anonymous):

Thanks

OpenStudy (karatechopper):

welcome:)